Elettra-Sincrotrone Trieste S.C.p.A. website uses session cookies which are required for users to navigate appropriately and safely. Session cookies created by the Elettra-Sincrotrone Trieste S.C.p.A. website navigation do not affect users' privacy during their browsing experience on our website, as they do not entail processing their personal identification data. Session cookies are not permanently stored and indeed are cancelled when the connection to the Elettra-Sincrotrone Trieste S.C.p.A. website is terminated.
More info
OK

HowTo&Troubleshooting

Hardware Troubleshooting

The Shutter is closed and it cannot be opened

You can check the Status of the FE valves here (BL BACH Monitor.VI):
or here: BACH BCS. Note that the valves are green when closed, red when open.

a) Try opening FE by the MainPanel.VI or the FE openclose.VI (how)..

b) Check if there is beam at the following link Machine Status: http://www.elettra.trieste.it/accelerator/machine_status.php

c) Check if the water fluxes of SM1, SM2, Entrance Slit at the very end of the beamline, after the monochromator, on the left are ok. If they are too low, adjust the levels (ask asap to your local contact to show you the position of these water gauges and flux regulator!).

In case the water flux is low, the corresponding element in the BACH BCS will be RED (Sph.Mirr.1,Sph.Mirr.2, Entr. Slit). If in normal operation they are GREEN.
See also: http://www.tasc.infm.it/research/bach/external.php#Troubleshooting

d) Check that all BL valves are open (in particular the first valve close to the front end)

e) Check the pressure of Sph.Mirr.1. The gauge is not working well and from time to time it needs to be restarted. Call the beamline scientists.

f) Check that the Gamma Monitor and that FE Hutch has been searched: BACS Supervisor Panel (Front-end hutch status,Stopper FE status,Gamma monitor,Ring shutter opening)

The FE is open, but there is no signal


      a) Check if there is beam at the following link Machine Status.

      b)Check that all BL valves are open (from the MainPanel.vi, press BL valves and then press the arrow to read the status)

      c)Check that the Machine shutter and the Machine valve are open (RED) (link: BACH BCS). These two elements are not controllable by the user. If you find them close you have to call the control room (8262 or 8263) to open them.

      d)Check that the Keithley (e.g. 25, reading Io) have been switched on using the VI Keithely Initialize 6430 inside the folder Counters. The first time initialise it at 10-6 (check first that entrance and exit slits are less than 50 micron). Then read Io using Counter read All.VI and select the more appropriate scale.


Io or Is are very noisy.


      Try to fix the preamplifier to the BNC(MHV) of the amnipulator using some aluminum foil in order to ensure a good electrical contact between the BNC of the manipulator and the preamp. Sometimes the weigh of the preamplifier losens the contact inside the MHV and small mechanical vibrations create electric noise. If this is not eough contact the BL scientist.

      Try to change the scale of the keithley. E.g. If you measure 3x10-9, set the scale to 10-8.

It is not possible to raise the energy above 0.6 keV in the Sputtering Power Controller.

      Solution: Check if the Extractor pin (n.2) is grounded (towards external shield of the ion gun). It should not be grounded! If a finite resistance is present, let some current flow trough this pin and ground until it is again isolated.

RECOVER FROM A POWER CUT

If there is a programmed powercut:

1) LIFT the MONO GRATING UP.
2) Switch OFF all the MULTIVACS of the ion pumps and gauges
3) Switch off all the motor controls
4) switch off all the PCs, including the MONO PC (ON1 and ON2, ask Roberto Borghes)
5) FE rack switch off by the the software instrument group (Billè, Turcinovich, Chenda, Borghes)

After the power recovery:
1) Check that all ION PUMPS are up and running. You might have to open the beamline racks.
2) Only if you are sure that the pressure id 10-8 or less, switch on the gauges.
3) MAIN CHAMBER: main turbo and turbo cart, Manip turbo cart and scroll; PREP: turbo+scroll
4) Check all valves along the beamline. Open them when p<10-9 mbar everywhere
5) In the FE racks: everything should be switched ON by the software instrument group (Billè, Turcinovich, Chenda, Borghes)

Run: MOTOR READ ALL.VI
Run: MOTOR RESTORE SPEED.VI

In the acquisition PC you will need to select again the MONO GRATING NUMBER and maybe to run again the MONO HOME procedure.

Some ref. values (June 24 2011):
PREHOR pitch=5606.9;Rll=0,x=-998.486(==enc);z=2.023
PREVER pitch=377.9 (enc=-9.365);Roll=-4.2,x=0.5;z=-0.235
ENT SLIT open=30 (enc 786.5); Y=-4.2
EX SLIT open=30 (enc=1048)
POSTHORA pitch=-46941.9;Rll=0,;z=-23.12
POSTHORB pitch=-67799.5;Rll=0,;z=-63

POSTVERA pitch=-8000.9;Roll=0.4
POSTVERA pitch=-6996.9;Roll=0.403
RIXS=-21.999

Heideneim controller switched off (power failure).

      Solution: Press ENT. Then move up and down the pin.

Seiko seiki STP300/400 controller (e.g. prep chamber) needs to be tuned (power failure).


.:. How to make the tuning adjustment of the STP300/400 SEIKO SEIKI PUMP to controller
The procedure (PDF file) can be found HERE.


 
Last Updated on Thursday, 13 September 2018 12:57